OpenStudy (anonymous):

\[\int_{-2}^{14} \frac{dx}{\sqrt[4]{x+2}}\]

OpenStudy (anonymous):

\[=\int_{-2}^{0} \frac{dx}{\sqrt[4]{x+2}}+\int_{0}^{14} \frac{dx}{\sqrt[4]{x+2}}\]

OpenStudy (anonymous):

^the interval is wrong

OpenStudy (anonymous):

improper integral here since the integrand is undefined at \(t=-2\)

OpenStudy (anonymous):